Calculate the molar mass of the following substances.
        Ethyne, C2H2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Ethyne, C2H2= 2 × 12 + 2 × 1 = 24 + 2 = 26 g
​​
Thanks